A track consists of a frictionless arc XY, which is a quarter circle of radius 2m and a rough horizontal section YZ.  Block A of mass (1kg) is released from rest at point X and collides instantaneously and inelastically with identical block B at point Y.  The two blocks move together to the right, maybe sliding past point P, which is a distance 2.5m from point Y.  The coefficient of kinetic friction between the blocks and the horizontal part of the track is 0.50.  Use g = 10m/s/s. How fast is block A traveling at point Y, just before it collides with block B?




1
Expert's answer
2021-03-09T07:47:14-0500

Apply energy conservation top to bottom-

 Potential energy = Kinetic energy

"\Rightarrow mgh=\dfrac{1}{2}mv^2"


"\Rightarrow(gR)=\dfrac{1}{2}v^2"


"v=\sqrt{2gR}\\

 =\sqrt{2\times 10\times 2}\\=\sqrt{40}\\=6.32\text{ms}^{-2}"